mykenzie:

Nick has 3 shirts: a white one, a black one, and a blue one. He also has two pairs of pants, one blue and one tan. What is the probability, if Nick gets dressed in the dark, that he winds up wearing the white shirt and tan pants?

Munkman450:

what are the answers?

mykenzie:

there r none

Munkman450:

o

mykenzie:

I need help finding the answer

Munkman450:

is there an answer u write in or click?

mykenzie:

I have to write

Munkman450:

o

mykenzie:

pls help

Munkman450:

K

Florisalreadytaken:

in the dark? haha thats the same supposing he is blind. so same with this one, two independent events occurring one after the other. lets assume that \( P(1) \) is the prob. for wearing a black shirt, whilst \( P(2) \) is the prob. for wearing tan pants can you determine the fractions?

mykenzie:

1/2 = 50%

Florisalreadytaken:

lets keep it as fractions there are 3 total shirts right? he will only wear one of them -- so we write: \[ P(1)= \frac{1}{3} \] as for the pants, we have got 2 possible options -- he also gets to wear only one pair of them : \[ P(2)=\frac{1}{2} \] \[ P(1 \cap 2) = P(1) \times P(2) \] \[ P(1 \cap 2) = \frac{1}{3} \times \frac{1}{2} \] \[ P(1 \cap 2) = ???? \]

mykenzie:

1/6

Florisalreadytaken:

yeah that looks right.

mykenzie:

k thxs